Suppose that 53% of the women who gave birth at a certain hospital last year were over 30 years old, and that 30% were unmarried. If 28% of the women were both unmarried and over 30, what is the probability that a woman who gave birth at the hospital was over 30 or unmarried (or both)?

Answer by stanbon(48569)

Suppose that 53% of the women who gave birth at a certain hospital last year were over 30 years old, and that 30% were unmarried. If 28% of the women were both unmarried and over 30, what is the probability that a woman who gave birth at the hospital was over 30 or unmarried (or both)?
-----------------------
General Formula: P(A or B) = P(A) + P(B) - P(A and B)
--------------------------------------------------------
P(>30 or unmarried) = P(>30) + P(unmarried) - P(>30 and unmarried)
P(>30 or unmarried) = 0.53 + 0.30 - 0.28 = 0.55
================================================
Cheers,
Stan H.